six years ago, rosa was four times old as her daughter. Ten years from now, she will be only twice as old. How old are they now???
Answer by mananth(16946)

daughter age = x
6 years ago she was x-6
Rosa's age 6 years ago= 4(x-6)
..
10 years later daughter = x+10
Rosa will be 4*(x-6)+16
..
4*(x-6)+16=2*(x+10)
4x-24+16=2X+20
4x-2x=24-16+20
2x=28
divide by 2
2x/2 = 28/2
x= 14 daughter's age now.
10 years later Rosa's age = 4(x-6)+16
plug value of x
4(14-6)+16
=32+16
=48
.
Now Rosa's age will be 48-10
= 38.